### v1

# New Algebraic Points on Curves

Abstract.
Let $C/\mathbb{Q}$ be a smooth projective absolutely irreducible curve of genus $\geq 2$. For an extension $L/\mathbb{Q}$, let $C(L)_{\mathrm{new}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:\;\mathbb{Q}(P)=L\}$. Let $n\geq 2$. We conjecture that $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of degree $n$ number fields $L$ when ordered by absolute discriminant. We prove our conjecture for $n=3$ with $C$ any of the modular curves $X_{0}(23)$, $X_{0}(29)$, $X_{0}(31)$ and $X_{0}(64)$. Moreover, we prove the analogue of this conjecture for the unit equation, again with $n=3$.
Key words and phrases:
hyperbolic curve, Diophantine stability, arithmetic statistics, modular curve
2020 Mathematics Subject Classification:
11G30, 11G18
1. Introduction
Let $C$ be a smooth projective and absolutely irreducible curve of genus $g$, defined over a number field $K$. Let $D$ be a reduced effective divisor on $C$ and consider the punctured curve X=CDX=C-D. The Euler characteristic of $X$ is given by $\chi(X)=2-2g-\deg(D)$. We say that $X$ is hyperbolic if $\chi(X)<0$. We let $\mathscr{X}/\mathcal{O}_{K}$ be a model for $X$. The famous Faltings–Siegel Theorem asserts that $\mathscr{X}(\mathcal{O}_{K})$ is finite. We mention three special cases.
  • If $D=0$ then $X=C$ is a complete curve, which is hyperbolic if and only if $g\geq 2$. In this case, $\mathscr{X}(\mathcal{O}_{K})=C(K)$ as $X$ is proper, and so Faltings–Siegel generalises Faltings’ theorem.
  • Let $X=\mathbb{P}^{1}-\{0,1,\infty\}$, which has Euler characteristic 1-1 and is therefore hyperbolic. It can be shown that $\varepsilon\in\mathscr{X}(\mathcal{O}_{K})$ if and only if $\varepsilon$ and $1-\varepsilon$ are units. Thus Faltings–Siegel generalises Siegel’s Theorem on the unit equation, i.e., the unit equation
    (1) $\varepsilon+\delta=1,\qquad\varepsilon,~\delta\in\mathcal{O}_{K}^{\times}$
    has finitely many solutions.
  • Let $E/K$ be an elliptic curve specified by an affine Weierstrass model
    $y^{2}+a_{1}xy+a_{3}y\;=\;x^{3}+a_{2}x^{2}+a_{4}x+a_{6},\qquad a_{i}\in\mathcal{O}_{K}$
    and let $\mathscr{O}$ denote the point at infinity. Let $X=E-\mathscr{O}$, which has Euler characteristic 1-1. Then the Faltings–Siegel theorem applied to $\mathscr{X}$ simply asserts that the affine model has finitely many points $(x,y)\in\mathcal{O}_{K}^{2}$; this is Siegel’s theorem for integral points on elliptic curves.
In the above cases, both the curve and the number field are fixed. This paper is motivated by the following question.
Question 1.
Given a hyperbolic curve $\mathscr{X}/K$ and a family $\mathcal{L}$ of number field extensions $L/K$, how does the set of integral points $\mathscr{X}(\mathcal{O}_{L})$ vary with the number field?
A typical instance of the above question is when $X=C$ is a complete curve of degree $g\geq 2$, and $\mathcal{L}$ is the set of all extensions $L/K$ of a fixed degree $n$. This instance of the question is well-considered in literature; for an excellent survey we recommend [25], but merely mention two famous results.
  • Merel’s celebrated Uniform Boundedness Theorem [16] asserts that all degree $n$ points on the modular curve $X_{1}(p)/\mathbb{Q}$ (with $p$ prime) are cuspidal, for $n<2\log_{3}(\sqrt{p}-1)$.
  • A theorem of Harris and Silverman [12] asserts that a (complete) curve $C/\mathbb{Q}$ of genus $\geq 2$ has infinitely many quadratic points if and only if $C$ is hyperelliptic or bielliptic. This theorem builds on a deep theorem of Faltings asserting that if $A/K$ is an abelian variety and $V$ is a subvariety not containing any abelian variety then $V(K)$ is finite.
In this paper we offer a conjectural statistical (partial) response to Question 1. Let $X/K$ be a hyperbolic curve as before, and let $L/K$ be an extension. The conjecture is concerned which what we call the set of $L$-new points that appear over $L$ on $X/K$; we define these by
$\mathscr{X}(\mathcal{O}_{L})_{\mathrm{new}}=\{P\in\mathscr{X}(\mathcal{O}_{L})\;:\;K(P)=L\}.$
We note that the $L$-new points are ones defined over $L$, but not over any strictly smaller extension. Before stating our conjecture, we explain what we mean by a family $\mathcal{L}$ of extensions $L/K$.
Definition 2.
Let $K$ be a number field. Let $n\geq 2$. Let $G_{1},\dotsc,G_{r}$ be transitive subgroups of $S_{n}$. Let $\upsilon_{1},\dotsc,\upsilon_{s}$ be a finite number of places of $K$, and let $L_{i}/K_{\upsilon_{i}}$ be degree $n$ étale algebras. By the degree $n$ family $\mathcal{L}$ with data $(G_{1},\dotsc,G_{r};L_{1},\dotsc,L_{s})$ we mean all extensions $L/K$ such that the following hold.
  • Write $\tilde{L}/K$ for the Galois closure of $L/K$. Then $\operatorname{Gal}(\tilde{L}/K)$ isomorphic to one of $G_{1},\dotsc,G_{s}$.
  • $L\otimes_{K}K_{\upsilon_{i}}\cong L_{i}$ for $i=1,\dotsc,r$.
We now state our conjecture, which we refer to as the statistical Diophantine stability conjecture for hyperbolic curves.
Conjecture 2.
Let $K$ be a number field and $\mathscr{X}/\mathcal{O}_{K}$ a model for a hyperbolic curve. Let $\mathcal{L}$ be a family of extensions $L/K$. Then $\mathscr{X}(\mathcal{O}_{L})_{\mathrm{new}}=\emptyset$ for 100% of $L\in\mathcal{L}$, when ordered by norms of discriminants.
If $X=C$ is a smooth projective absolutely irreducible curve over a number field $K$, then
$\mathscr{X}(\mathcal{O}_{L})_{\mathrm{new}}\;=\;C(L)_{\mathrm{new}}\;\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:K(P)=L\}.$
Thus Conjecture 2 becomes the following simpler statement.
Conjecture 3.
Let $K$ be a number field and let $C/K$ be a smooth projective absolutely irreducible curve of genus $\geq 2$. Let $\mathcal{L}$ be a family of extensions $L/K$. Then $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of $L\in\mathcal{L}$, when ordered by norms of discriminants.
Remarks.
  • The idea of Diophantine stability first appears in a different context due to Mazur and Rubin [15].
  • It is tempting to conjecture that $\mathscr{X}(\mathcal{O}_{L})=\mathscr{X}(\mathcal{O}_{K})$ for $100\%$ of $L\in\mathcal{L}$. This however is false. For example, let $\mathcal{L}$ be all quartic number fields $L/\mathbb{Q}$. Let $X/\mathbb{Q}$ be a hyperbolic curve such that $\mathscr{X}(\mathbb{Z})=\emptyset$, but $\mathscr{X}(\mathbb{Z}[(1+\sqrt{5})/2])\neq\emptyset$. A positive proportion of quartic number fields $L$, when ordered by absolute discriminant, contain $\mathbb{Q}(\sqrt{5})$ ([4, Corollary 1.2]) and therefore $\mathscr{X}(\mathcal{O}_{L})\neq\mathscr{X}(\mathbb{Z})$ for those fields.
We mention a few special cases of Conjecture 2 in the literature.
  1. (I)
    Let $n$ be a positive integer with $3\nmid n$. Let $\mathcal{L}$ be the degree $n$ number fields $L/\mathbb{Q}$ where 33 splits completely. A theorem of Triantafillou [24] asserts that $(\mathbb{P}^{1}-\{0,1,\infty\})(\mathcal{O}_{L})=\emptyset$ for $L\in\mathcal{L}$. Similar theorems for many other “locally restricted” families $\mathcal{L}$ were proved by Freitas, Kraus and Siksek [8, 9, 10, 11].
  2. (II)
    Let $\ell$ be a prime and let $\mathcal{L}$ be the set of cyclic degree $\ell$ number fields $L/\mathbb{Q}$. Let $C/\mathbb{Q}$ be a curve and let $P\in C(\mathbb{Q})$ be a rational point. Write $J_{C}$ for the Jacobian of $C$ and suppose that $J_{C}(\mathbb{Q})=0$. Let $X=C-\{P\}$. Siksek [22] shows that $\mathscr{X}(\mathcal{O}_{L})=\emptyset$ for $100\%$ of $L\in\mathcal{L}$, subject to a mild restriction on the mod $\ell$ representation of $J_{C}$.
The purpose of this paper is to prove Conjecture 2 for several hyperbolic curves, with $\mathcal{L}$ the set of all cubic fields. All these hyperbolic curves have infinitely many cubic points.
Theorem 4.
$(\mathbb{P}^{1}-\{0,1,\infty\})(\mathcal{O}_{L})=\emptyset$ for $100\%$ of cubic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
Theorem 5.
Let $X=X_{0}(23)$, $X_{0}(29),X_{0}(31)$ or $X_{0}(64)$ (these have genera $2$, 22, 22, 33 respectively). Then $X(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of cubic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
To motivate what comes next, we mention the following “single source theorem” due to the authors [14, Theorem 1.1]. In this context, by a primitive degree $n$ point on $C/K$, we mean an algebraic point $P\in C(\overline{K})$ such that $K(P)/K$ has degree $n$, and the Galois group of its Galois closure is a primitive subgroup of $S_{n}$.
Theorem.
Let $K$ be a number field. Let $C/K$ be a curve of genus $g$, and write $J$ for the Jacobian of $C$. Let $n\geq 2$ and suppose
(3) $\begin{cases}g>(n-1)^{2}&\text{if $n\geq 3$}\\ g\geq 3&\text{if $n=2$}.\end{cases}$
Suppose that $A(K)$ is finite for every abelian subvariety $A/K$ of $J$ of dimension $\leq n/2$. If $C$ has infinitely many primitive points of degree $n$, then there is a degree $n$ morphism $\varphi:C\rightarrow\mathbb{P}^{1}$ defined over $K$ such that all but finitely many primitive degree $n$ divisors are fibres $\varphi^{*}(t)$ with $t\in\mathbb{P}^{1}(K)$.
This theorem is called the “Single Source Theorem” since, with finitely many exceptions, all primitive degree $n$ points come from a single source which is the morphism $\varphi:C\rightarrow\mathbb{P}^{1}$. In view of our Conjecture 2, it makes sense to consider the family of degree $n$ number fields we obtain from the fibres of this morphism, and to attempt to show that this family contains $0\%$ of degree $n$ number fields. We shall give a result in this spirit.
Let $C$ be a smooth projective absolutely irreducible curve over $K$, and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a morphism of degree $n\geq 2$. Then the curve $C$ possesses a plane affine (possibly singular) model
C:F(T,Y)=0C\;:\;F(T,Y)=0
where $F(T,Y)\in\mathbb{Q}(T)[Y]$ has degree $n$ in $Y$, is irreducible over $\overline{\mathbb{Q}}(T)$, and where the morphism $\varphi$ is given by the map $(T,Y)\mapsto T$. Hilbert’s irreducibility theorem asserts the existence of a thin set $\Theta\subset\mathbb{Q}$, such that for $t\in\mathbb{Q}-\Theta$, the polynomial $F(t,Y)\in\mathbb{Q}[Y]$ is irreducible of degree $n$. Thus $C$ has infinitely many degree $n$ points. For $t\in\mathbb{Q}-\Theta$ we write $K_{t}$ for the degree $n$ number field defined by $F(t,Y)=0$, and let $\Delta(\mathcal{O}_{K_{t}})$ be the discriminant of the ring of integers $\mathcal{O}_{K_{t}}$ of $K_{t}$. We define the discriminant set of $F$ to be
$\operatorname{Disc}_{F}\;=\;\{\lvert\Delta(\mathcal{O}_{K_{t}})\rvert\;:\;t\in\mathbb{Q}-\Theta\}.$
We shall give a mild sufficient condition for the set $\operatorname{Disc}_{F}$ to have natural density 0 among the positive integers. Whilst this is a long way off from Conjecture 2, it does harmonise with it.
Theorem 6.
Let $\Delta_{Y}(F)\in\mathbb{Q}(T)$ be the discriminant of $F$ with respect to $Y$, and write this as
$\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
where $g\in\mathbb{Q}[T]$ is a squarefree polynomial and $h(T)\in\mathbb{Q}(T)$ is a rational function. Suppose $g$ has even degree. Let $G=\operatorname{Gal}(g)$. Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Then there is some $0<\alpha<1$ such that, as $X\rightarrow\infty$,
$(\operatorname{Disc}_{F}\cap[1,X])\;=\;O\left(\frac{X}{(\log{X})^{\alpha}}\right).$
In particular, the discriminant set $\operatorname{Disc}_{F}$ has density 0 among the natural numbers.
We are grateful to Maarten Derickx and Victor Flynn for helpful discussions.
2. Counting integers with certain restricted factorisations
Let $\mathbb{P}$ be the set of prime numbers and let $\mathcal{P}\subseteq\mathbb{P}$. Following Serre [21], we call $\mathcal{P}$ regular of density $\alpha>0$ if
(4) $\sum_{p\in\mathcal{P}}\frac{1}{p^{s}}\;=\;\alpha\cdot\log\left(\frac{1}{s-1}\right)\;+\;\theta_{A}(s)$
where $\theta_{A}$ extends to a holomorphic function on $\operatorname{Re}(s)\geq 1$. We call the set $\mathcal{P}$ Frobenian of density $\alpha>0$ if there exists a finite Galois extension $L/\mathbb{Q}$ and a subset $\mathscr{C}$ of $G=\operatorname{Gal}(L/\mathbb{Q})$, such that
  • $\mathscr{C}$ is a union of conjugacy classes in $G$;
  • $\alpha=\#\mathscr{C}/\#G$;
  • for every sufficiently large prime $p$, we have $p\in\mathcal{P}$ if and only if $\sigma_{p}\in\mathscr{C}$ where $\sigma_{p}$ is a Frobenius element of $G$ corresponding to $p$.
By the Chebotarev Density Theorem [21, Proposition 1.5], if $\mathcal{P}$ is Frobenian of density $\alpha>0$ then it is regular of density $\alpha>0$.
Theorem 7.
Let $S$ be a set of primes of regular density $0<\alpha<1$. Let $N_{S}(X)$ be the number of positive integers $n\leq X$ such that $\operatorname{ord}_{p}(n)$ is even for all $p\in S$. Then, there exists some constant $C_{S}>0$ such that
$N_{S}(X)\;\thicksim\;C_{S}\cdot\frac{X}{(\log{X})^{\alpha}}.$
Proof.
Let
$a_{n}=\begin{cases}1&\text{if $\operatorname{ord}_{p}(n)$ is even for all $p\in S$}\\ 0&\text{otherwise}.\end{cases}$
Then $N_{S}(X)=\sum_{n\leq X}a_{n}$. We shall make use of a Tauberian theorem due to Delange [23, page 350] to estimate $N_{S}(X)$.
Consider the Dirichlet series
$D(s)=\sum_{n=1}^{\infty}\frac{a_{n}}{n^{s}},$
which defines a holomorphic function on $\operatorname{Re}(s)>1$. We need to analyse the behaviour of $D(s)$ in the neighbourhood of $s=1$. For $\operatorname{Re}(s)>1$ we write $D(s)$ as an Euler product
$\begin{split}D(s)&=\prod_{p\in S}\left(1+\frac{1}{p^{2s}}+\frac{1}{p^{4s}}+\cdots\right)\cdot\prod_{p\notin S}\left(1+\frac{1}{p^{s}}+\frac{1}{p^{2s}}+\cdots\right)\\ &=\prod_{p\in S}\left(1-\frac{1}{p^{2s}}\right)^{-1}\cdot\prod_{p\notin S}\left(1-\frac{1}{p^{s}}\right)^{-1}.\end{split}$
Thus
$\log(D(s))\;=\;\sum_{p\notin S}\frac{1}{p^{s}}\;+\;\theta(s)$
where $\theta$ is holomorphic at $s=1$. Since $\mathbb{P}-S$ has regular density $1-\alpha>0$, we conclude from (4) that
$\log(D(s))\;=\;(1-\alpha)\cdot\log\left(\frac{1}{s-1}\right)+\phi(s)$
where $\phi$ is holomorphic at $s=1$. Hence
$D(s)\;=\;\frac{g_{0}(s)}{(s-1)^{1+\omega}}.$
where $g_{0}(s)=\exp(\phi(s))$ is holomorphic and non-vanishing at $s=1$, and $\omega=-\alpha$; the notation here is chosen to match that of Theorem II.7.28 of [23]. We note that $1<\omega<0$. Applying the aforementioned theorem gives
$\sum_{n\leq X}a_{n}\;\thicksim\;\frac{g_{0}(1)}{\Gamma(\omega+1)}\cdot X(\log{X})^{\omega}.$
This gives the theorem with $C_{S}=\exp(\phi(1))/\Gamma(1-\alpha)$. ∎
3. Proof of Theorem 6
Lemma 8.
Suppose $F$ satisfies the hypotheses of Theorem 6. Then there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even.
Proof.
Let $t\in\mathbb{Q}-\Theta$, and write $t=a/b$ where $a$, $b$ are coprime integers. We note that
$\Delta(\mathcal{O}_{K_{t}})/\Delta_{Y}(F(t,Y))\;\in\;(\mathbb{Q}^{\times})^{2}.$
Thus
$\Delta(\mathcal{O}_{K_{t}})/g(t)\;\in\;(\mathbb{Q}^{\times})^{2}.$
As $g$ has even degree, say $m$, we can write
$g(t)\;=\;\frac{G(a,b)}{b^{m}}$
where $G$ is homogeneous, and so
$\Delta(\mathcal{O}_{K_{t}})/G(a,b)\;\in\;(\mathbb{Q}^{\times})^{2}.$
Thus, for any prime $p$,
$\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))\;\equiv\;\operatorname{ord}_{p}(G(a,b))\pmod{2}$.
By assumption, there is some $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Therefore, by the Chebotarev Density Theorem, there is a set of primes $S$, having positive regular density, such that the congruence $G(u,v)\equiv 0\pmod{p}$ does not have solutions with coprime $u$, $v$. Hence for $p\in S$, we have $\operatorname{ord}_{p}(G(a,b))=0$, so $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even. ∎
Proof of Theorem 6.
Theorem 6 follows immediately from Theorem 7 and Lemma 8. ∎
4. Counting low degree fields with local restrictions
In this section we recount some well-known theorems which count number fields of degrees $2$, 33, 44, 55, with certain local restrictions. We shall not need the most general form of these theorems, and content ourselves with stating simplified special cases. The first concerns quadratic fields, and is a special case of [7, Proposition A.1].
Theorem 9 (Ellenberg, Pierce and Wood).
Let $T$ be a finite set of primes. Write $N_{T}^{\pm}(X)$ for the number of isomorphism classes of quadratic fields $K$ satisfying
  • $0<\pm\Delta_{K}\leq X$;
  • all $p\in T$ are unramified in $K$.
Then
$N_{T}^{\pm}(X)\;=\;C_{T}\cdot X+O_{T}(X^{1/2})$
where $C=1/2\zeta(2)$ and
$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{1}{p+1}\right).$
Most importantly for us, we shall need a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications. The following two corollaries are easy consequences of that theorem.
Corollary 10.
Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $K$ satisfying
  1. (i)
    $\lvert\Delta_{K}\rvert\leq X$;
  2. (ii)
    $\operatorname{ord}_{p}(\Delta_{K})=0$ for all $p\in T$.
Then
$N_{T}(X)\;=\;C_{T}\cdot X+O_{T}(X^{5/6})$
where
$C_{T}\;=\;\frac{1}{3\zeta(3)}\cdot\prod_{p\in T}\left(1-\frac{p+1}{p^{2}+p+1}\right).$
Corollary 11.
Let $T$ be a finite set of primes, with $3\notin T$. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $K$ satisfying
  1. (i)
    $\lvert\Delta_{K}\rvert\leq X$;
  2. (ii)
    $\operatorname{ord}_{p}(\Delta_{K})$ is even for all $p\in T$.
Then
$N_{T}(X)\;=\;C_{T}\cdot X+O_{T}(X^{5/6})$
where
$C_{T}\;=\;\frac{1}{3\zeta(3)}\cdot\prod_{p\in T}\left(1-\frac{p}{p^{2}+p+1}\right).$
Proof of Corollaries 10 and 11.
Theorem 1.3 of [1] expresses the density $C_{T}^{\pm}$ as a product of local densities, which depend on the splitting type of $p$ in $K$ (either totally split, partially split, inert, partially ramified, totally ramified). In Corollary 10 the allowed splitting types for $p\in T$ are totally split, partially split, inert, and the corollary follows immediately from the aforementioned [1, Theorem 1.3].
Next, we note that for $p\neq 3$, the valuation $\operatorname{ord}_{p}(\Delta_{K})$ is either 0, 11 or $2$ according to whether $p$ is unramified in $K$, partially ramified in $K$, or totally ramified in $K$. Thus, condition (ii) is equivalent to saying that for every $p\in T$ the splitting type of $p$ is either totally split, partially split, inert or totally ramified. The result again follows from [1, Theorem 1.3]. ∎
The following is a special case of [7, Theorem 4.1].
Theorem 12 (Ellenberg, Pierce and Wood).
Let $T$ be a finite set of primes. Write $\tilde{N}_{T}(X)$ for the number of isomorphism classes of non-$D_{4}$ quartic fields $K$ satisfying
  • $0<\lvert\Delta_{K}\rvert\leq X$;
  • all $p\in T$ are unramified in $K$.
Then
$\tilde{N}_{T}(X)\;=\;C_{T}\cdot X+O_{T,\epsilon}(X^{23/24+\epsilon})$
where
$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{p^{2}+2p+1}{p^{3}+p^{2}+2p+1}\right).$
and $C>0$ is some fixed absolute constant.
Cohen, Diaz y Diaz and Olivier [4] show that the number of $D_{4}$-extensions with absolute discriminant at most $X$ is $\tilde{c}\cdot X$ where $c\approx 0.052326$, but there does not seem to be a count for $D_{4}$-extension with local restrictions.
Finally, the following is a special case of [7, Theorem 5.1].
Theorem 13 (Ellenberg, Pierce and Wood).
Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of quintic fields $K$ satisfying
  • $0<\lvert\Delta_{K}\rvert\leq X$;
  • all $p\in T$ are unramified in $K$.
Then
$N_{T}(X)\;=\;C_{T}\cdot X+O_{T,\epsilon}(X^{199/200+\epsilon})$
where
$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{p^{3}+2p^{2}+2p+1}{p^{4}+p^{3}+2p^{2}+2p+1}\right).$
and $C>0$ is some fixed absolute constant.
Lemma 14.
Let $R=\{p_{1},p_{2},p_{3},\dotsc\}$ be a set of primes having regular density >0>0. Let $f$, $g$ be monic polynomials with coefficients in $\mathbb{R}$, with $\deg(g)=\deg(f)+1$, and suppose $g(p_{i})\neq 0$ for all $i$. Then
$\lim_{k\rightarrow\infty}\prod_{i=1}^{k}\left(1-\frac{f(p_{i})}{g(p_{i})}\right)=0.$
Proof.
It is sufficient to show that $\sum_{p\in R}f(p)/g(p)$ diverges. Suppose $\sum_{p\in R}f(p)/g(p)$ converges. We note that $xf(x)$ and $g(x)$ are monic polynomials of the same degree, therefore $g(p)/pf(p)$ is bounded, say by $B>0$, on $R$. We note that $1/p\leq B\cdot f(p)/g(p)$ for $p\in R$. Thus $\sum_{p\in R}1/p$ converges. Let $\delta$ be the regular density of $R$, which by assumption is positive. From (4) we have
$\delta=\lim_{\sigma\rightarrow 1^{+}}\frac{\sum_{p\in R}\frac{1}{p^{\sigma}}}{-\log(\sigma-1)}.$
However, as $\sigma\rightarrow 1^{+}$,
$\frac{\sum_{p\in R}\frac{1}{p^{\sigma}}}{-\log(\sigma-1)}\;\leq\;\frac{\sum_{p\in R}\frac{1}{p}}{-\log(\sigma-1)}\rightarrow 0$
since $\sum_{p\in\mathbb{P}}1/p^{\sigma}=-\log(\sigma-1)+O(1)$. This contradicts the assumption that $\delta>0$. ∎
Corollary 15.
Let $S$ be a set of primes having regular density <1<1.
  1. (a)
    Let $d=2$, 33 or $5$. Then the proportion of degree $d$ number fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
  2. (b)
    Let $d=4$. Then the proportion of non-$D_{4}$ quartic fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
Proof.
Write $\mathcal{F}(X)$ for the set of quintic fields with absolute discriminant at most $X$, and write $\mathcal{F}_{S}(X)$ for the subset whose discriminants are supported on $S$. To prove the corollary for $d=5$, we want to show that
(5) $\lim_{X\rightarrow\infty}\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}=0.$
Let $p_{1},\dotsc,p_{k}$ be primes belonging to $R:=\mathbb{P}-S$. Then
$\mathcal{F}_{S}(X)\;\subset\;\{K\;:\;\text{$K\in\mathcal{F}(X)$, $p_{1},\dotsc,p_{k}$ unramified in $K$}\}.$
Hence
$0\leq\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}\leq\frac{N_{p_{1},\dotsc,p_{k}}(X)}{N_{\emptyset}(X)}\rightarrow\prod_{i=1}^{k}\left(1-\frac{p_{i}^{3}+2p_{i}^{2}+2p_{i}+1}{p_{i}^{4}+p_{i}^{3}+2p_{i}^{2}+2p_{i}+1}\right)$
as $X\rightarrow\infty$, where the limit follows from Theorem 13. From Lemma 14, by choosing $p_{1},\dotsc,p_{k}$ be a suitable (finite) subset of $R$, we can ensure that the product on the right hand-side is as small as we like. This completes the proof of (a) for $d=5$. The other cases are similar. ∎
Corollary 16.
Let $S$ be a set of primes having regular density >0>0. Then the proportion of cubic fields with discriminants having even valuation at all $p\in S$ is 0.
Proof.
Write $\mathcal{F}(X)$ for the set of cubic fields with absolute discriminant at most $X$, and write $\mathcal{F}_{S}(X)$ for the subset whose discriminants have even valuation at all $p\in S$. Again we want to prove (5). Let $p_{1},\dotsc,p_{k}$ be primes belonging to $S$. Then
$\mathcal{F}_{S}(X)\;\subset\;\{K\;:\;\text{$K\in\mathcal{F}(X)$, where $\operatorname{ord}_{p_{i}}(\Delta_{K})$ is even for all $i=1,\dotsc,k$}\}.$
Hence
$0\leq\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}\leq\frac{N_{p_{1},\dotsc,p_{k}}(X)}{N_{\emptyset}(X)}\rightarrow\prod_{i=1}^{k}\left(1-\frac{p_{i}}{p_{i}^{2}+p_{i}+1}\right)$
as $X\rightarrow\infty$, where the limit follows from Corollary 11. From Lemma 14, by choosing $p_{1},\dotsc,p_{k}$ be a suitable subset of $S$, we can ensure that the product on the right hand-side is as small as we like. This completes the proof. ∎
Corollary 17.
Let $F(T,Y)\in\mathbb{Q}(T)[Y]$ has degree $3$ in $Y$, and be irreducible over $\mathbb{Q}(T)$, and $\Theta$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ either has degree <n<n or is reducible (this set is necessarily thin by Hilbert’s irreducibility theorem). Let $\Delta_{Y}(F)\in\mathbb{Q}(T)$ be the discriminant of $F$ with respect to $Y$, and write this as
$\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
where $g\in\mathbb{Q}[T]$ is a squarefree polynomial and $h(T)\in\mathbb{Q}(T)$ is a rational function. Suppose $g$ has even degree. Let $G=\operatorname{Gal}(g)$. Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Let $\mathcal{F}^{\prime}$ be the set of cubic fields we obtain from $F(t,Y)=0$ with $t\in\mathbb{Q}-\Theta$. Then $100\%$ of cubic fields, ordered by discriminant, do not belong to the family $\mathcal{F}^{\prime}$.
Proof.
By Lemma 8, there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even, for any $K_{t}\in\mathcal{F}^{\prime}$. The desired result follows from Corollary 16. ∎
5. Proof of Theorem 4
Let $L$ be a number field. Nagell [19] calls a unit $\varepsilon\in\mathcal{O}_{L}^{\times}$ exceptional if $1-\varepsilon\in\mathcal{O}_{L}^{\times}$. The number field $L$ is called exceptional if it possesses an exceptional unit. Thus $\varepsilon$ is exceptional if and only if $(\varepsilon,1-\varepsilon)$ is a solution to the unit equation (1), and $L$ is exceptional if and only if the unit equation has solutions in $L$. Nagell [17], [18], has determined all exceptional number fields where the unit rank is 0 or $1$. Nagell finds [18, Section 2] that the only exceptional quadratic fields are $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-3})$ which contain exceptional units $(1+\sqrt{5})/2$ and $(1-\sqrt{-3})/2$ respectively, and the only exceptional complex cubic fields are the ones with discriminants 23-23 and 31-31. He also showed [18, Sections 3–5] that the only exceptional real cubic fields are of the form $\mathbb{Q}(\varepsilon)$ where $\varepsilon$ is a root of
$f_{t}(X)=X^{3}+(t-1)X^{2}-tX-1,\qquad t\in\mathbb{Z},\quad t\geq 3$
or of
$g_{t}(X)=X^{3}+tX^{2}-(t+3)X+1,\qquad t\in\mathbb{Z},\quad t\geq-1,$
and in both cases $\varepsilon$ is an exceptional unit. Whilst the number of exceptional cubic fields is infinite, we are able to show the following.
Theorem 18.
$100\%$ of cubic number fields are non-exceptional, when ordered by absolute discriminant.
This theorem is of course equivalent to Theorem 4.
Proof.
By the aforementioned work of Nagell [18], it remains to prove that $100\%$ of real cubic fields are non-exceptional, when ordered by absolute discriminant. We will first consider the case that $\varepsilon$ is an exceptional cubic unit and a root of the polynomial
$g_{t}(X)=X^{3}+tX^{2}-(t+3)X+1$
for some $t\in\mathbb{Z}_{\geq-1}$. The discriminant of $g_{t}(X)$ is $(t^{2}+3t+9)^{2}$, and therefore $\mathbb{Q}(\varepsilon)$ has Galois group $C_{3}$. This completes the proof in this case as Cohn and Davenport–Heilbronn have shown that $100\%$ of cubic fields have Galois group $S_{3}$, when ordered by absolute discriminant [5, 6]. Now suppose $\varepsilon$ is an exceptional cubic unit and a root of the polynomial
$f_{t}(X)=X^{3}+(t-1)X^{2}-tX-1$
for some $t\in\mathbb{Z}_{\geq 3}$. The discriminant of $f_{t}(X)$, with respect to $X$, is given by
$F(t)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Delta_{X}(f_{t})=t^{4}+6t^{3}+7t^{2}-6t-31.$
It follows that if a prime $p$ divides the discriminant of $\mathcal{O}_{\mathbb{Q}(\varepsilon)}$ then it divides $t^{4}+6t^{3}+7t^{2}-6t-31$ and hence the polynomial $F(X)$ has a root modulo $p$. Write $S$ for the set of primes $p$ such that $F$ has a root modulo $p$. The Galois group of $F$ is $D_{4}$ and an easy application of the Chebotarev Density Theorem shows that $S$ has Frobenian (and hence regular) density $3/8$.
Write $\mathcal{F}_{S}$ for the set of cubic fields with discriminants supported on $S$. By the above, the exceptional fields with Galois group $S_{3}$ all belong to $\mathcal{F}_{S}$. The latter has density 0 by Corollary 15. This completes the proof.
6. Proof of Theorem 5
We will give the details for $X_{0}(64)$, which has genus $3$. The other cases are similar. The paper of Ozman and Siksek [20] gives the following convenient model for $X_{0}(64)$ as a plane quartic
$X_{0}(64)\;:\;x^{3}z+4xz^{3}-y^{4}\;=\;0.$
We point in passing that this is isomorphic to the Fermat quartic, and that Bremner and Choudhry [3, Theorem 3.1] have shown that all cubic points on this curve have Galois group $S_{3}$. We shall parametrize the cubic points. We note the following four rational points on the above model:
$P_{1}=[0:0:1],\;P_{2}=[2:2:1],\;P_{3}:=[2:-2:1],\;P_{4}:=[1:0:0].$
Let
$D_{1}=-P_{1}+3P_{2}-3P_{3}+P_{4},\;D_{2}=-P_{1}+2P_{2}-2P_{3}+P_{4},\;D_{3}=2P_{1}-2P_{2}+P_{3}-P_{4}.$
Using the Magma programs accompanying the paper [20] we find that
$J(\mathbb{Q})=(\mathbb{Z}/2\mathbb{Z})\cdot[D_{1}]+(\mathbb{Z}/4\mathbb{Z})\cdot[D_{2}]+(\mathbb{Z}/4\mathbb{Z})\cdot[D_{3}].$
Now let $Q$ be a degree $3$ point on $X_{0}(64)$, and let $D$ be the corresponding irreducible effective degree $3$ divisor. Then $D-3P_{1}$ is a degree 0 divisor, and so $D-3P_{1}]\in J(\mathbb{Q})$. Thus,
$D\sim\underbrace{3P_{1}+aD_{1}+bD_{2}+cD_{3}}_{D_{a,b,c}},\qquad 0\leq a\leq 1,\quad-1\leq b,c\leq 2.$
Using Magma [2] we computed the Riemann–Roch space $L(D_{a,b,c})$ for the $32$ possible combinations $(a,b,c)$. For all but $4$ of these, the Riemann–Roch dimension is $1$, and therefore $D=D_{a,b,c}$. For the remaining $4$, namely $(a,b,c)=(0,0,0)$, $(0,1,0)$, $(0,1,1)$, $(1,2,1)$, the Riemann–Roch dimension is $2$, and therefore there are infinitely many effective degree $3$ divisors linearly equivalent to $D_{a,b,c}$. Consider the case $(a,b,c)=(0,0,0)$. Then $D_{(0,0,0)}=3P_{1}$, and a basis for $L(D_{(0,0,0)})$ is given by $1$, $\varphi=y/x$, where these are considered as elements of the function field $\mathbb{Q}(X_{0}(64))$. Thus
$D=\operatorname{div}(\varphi-t)+3P_{1}$
for some $t\in\mathbb{Q}$. We can also consider $\varphi$ as a degree $3$ morphism $X_{0}(64)\rightarrow\mathbb{P}^{1}$, and then $D=\varphi^{-1}(t)$. This gives an infinite family of effective degree $3$ divisors on $X_{0}(64)$. More concretely, write $T=y/z$, and $Y=x/z$ and let
$F(T,Y)=4Y^{3}+Y-T^{4};$
then $F(T,Y)=0$ is an affine plane model for $X_{0}(64)$ where the morphism $\varphi$ is simply given by $(T,Y)\mapsto T$. In the notation of Corollary 17, we find
$\Delta_{Y}(F)=-432T^{8}-16=g(T)\cdot h(T)^{2},\qquad g(T)=-27T^{8}-1,\quad h(T)=4.$
As $g$ is irreducible, the Galois group will contain elements that act freely on the roots by a theorem of Jordan [13]. More precisely, the Galois group is $\operatorname{Gal}(g)=D_{8}\rtimes C_{2}$ which has order $32$, and $21$ of the elements act freely on the roots. By Corollary 17, $100\%$ of cubic fields will not arise from equations of the form $F(t,Y)=0$ with $t\in\mathbb{Q}$. The other three cases $(a,b,c)=(0,1,0)$, $(0,1,1)$, $(1,2,1)$ are similar. We refer the reader to
https://github.com/MaleehaKhawaja/NewPoints
for the Magma code supporting the proof of this theorem.

### v2

@@
 Let $C/\mathbb{Q}$ be a smooth projective absolutely irreducible curve of genus $\geq 2$.
-For an extension $L/\mathbb{Q}$, let $C(L)_{\mathrm{new}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:\;\mathbb{Q}(P)=L\}$.
+For an extension $L/\mathbb{Q}$, let $C(L)_{\mathrm{new}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:\;\mathbb{Q}(P)=L\}$; this is the set of points on $C$ defined over $L$, but not over any strictly smaller field.
 Let $n\geq 2$.
-We conjecture that $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of degree $n$ number fields $L$ when ordered by absolute discriminant.
-We prove our conjecture for $n=3$ with $C$ any of the modular curves $X_{0}(23)$, $X_{0}(29)$, $X_{0}(31)$ and $X_{0}(64)$.
-Moreover, we prove the analogue of this conjecture for the unit equation, again with $n=3$.
+We conjecture that $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of degree $n$ number fields $L$, when ordered by absolute discriminant.
+For degrees $n=2$, 33, we give sufficient criteria for our conjecture to hold in terms of an explicit model for $C$.
+For general $n$ we prove a theorem that harmonises with the conjecture.
+In particular, we verify our conjecture in these cases
+  • $n=2$; $C=X_{0}(N)$ for the $18$ values $N\neq 37$ such that $X_{0}(N)$ is hyperelliptic.
+  • $n=3$; $C=X_{0}(23)$, $X_{0}(29)$, $X_{0}(31)$, $X_{0}(64)$.
+Moreover, we prove the analogue of our conjecture for the unit equation, again with $n=3$.
 Key words and phrases:
@@
 We say that $X$ is hyperbolic if $\chi(X)<0$.
-We let $\mathscr{X}/\mathcal{O}_{K}$ be a model for $X$.
+We write $\mathcal{O}_{K}$ for the ring of integers of $K$, and we let $\mathscr{X}/\mathcal{O}_{K}$ be a model for $X$ over $\mathcal{O}_{K}$.
 The famous Faltings–Siegel Theorem asserts that $\mathscr{X}(\mathcal{O}_{K})$ is finite.
@@
 It can be shown that $\varepsilon\in\mathscr{X}(\mathcal{O}_{K})$ if and only if $\varepsilon$ and $1-\varepsilon$ are units.
-Thus Faltings–Siegel generalises Siegel’s Theorem on the unit equation, i.e., the unit equation
+Thus Faltings–Siegel generalises Siegel’s Theorem which states that the unit equation
     (1) $\varepsilon+\delta=1,\qquad\varepsilon,~\delta\in\mathcal{O}_{K}^{\times}$
@@
 Given a hyperbolic curve $\mathscr{X}/K$ and a family $\mathcal{L}$ of number field extensions $L/K$, how does the set of integral points $\mathscr{X}(\mathcal{O}_{L})$ vary with the number field?
-A typical instance of the above question is when $X=C$ is a complete curve of degree $g\geq 2$, and $\mathcal{L}$ is the set of all extensions $L/K$ of a fixed degree $n$.
-This instance of the question is well-considered in literature; for an excellent survey we recommend [25], but merely mention two famous results.
+A typical instance of the above question is when $X=C$ is a complete curve of genus $g\geq 2$, and $\mathcal{L}$ is the set of all extensions $L/K$ of a fixed degree $n$.
+This instance of the question is well-considered in the literature; for an excellent survey we recommend [35], but merely mention two famous results.
   • Merel’s celebrated Uniform Boundedness Theorem [21] asserts that all degree $n$ points on the modular curve $X_{1}(p)/\mathbb{Q}$ (with $p$ prime) are cuspidal, for $n<2\log_{3}(\sqrt{p}-1)$.
-  • A theorem of Harris and Silverman [12] asserts that a (complete) curve $C/\mathbb{Q}$ of genus $\geq 2$ has infinitely many quadratic points if and only if $C$ is hyperelliptic or bielliptic.
+  • A theorem of Harris and Silverman [15] asserts that if a (complete) curve $C/\mathbb{Q}$ of genus $\geq 2$ has infinitely many quadratic points then $C$ is hyperelliptic or bielliptic.
 This theorem builds on a deep theorem of Faltings asserting that if $A/K$ is an abelian variety and $V$ is a subvariety not containing any abelian variety then $V(K)$ is finite.
-In this paper we offer a conjectural statistical (partial) response to Question 1.
+In this paper we offer a conjectural statistical (partial) response to Question 1; this is Conjecture 2 stated below.
 Let $X/K$ be a hyperbolic curve as before, and let $L/K$ be an extension.
-The conjecture is concerned which what we call the set of $L$-new points that appear over $L$ on $X/K$; we define these by
+The conjecture is concerned with what we call the set of $L$-new points that appear over $L$ on $X/K$; we define these by
 $\mathscr{X}(\mathcal{O}_{L})_{\mathrm{new}}=\{P\in\mathscr{X}(\mathcal{O}_{L})\;:\;K(P)=L\}.$
-We note that the $L$-new points are ones defined over $L$, but not over any strictly smaller extension.
+We note that the $L$-new points are ones defined over $L$, but not over any strictly smaller extension of $K$.
 Before stating our conjecture, we explain what we mean by a family $\mathcal{L}$ of extensions $L/K$.
@@
 Let $\upsilon_{1},\dotsc,\upsilon_{s}$ be a finite number of places of $K$, and let $L_{i}/K_{\upsilon_{i}}$ be degree $n$ étale algebras.
-By the degree $n$ family $\mathcal{L}$ with data $(G_{1},\dotsc,G_{r};L_{1},\dotsc,L_{s})$ we mean all extensions $L/K$ such that the following hold.
+By the degree $n$ family $\mathcal{L}$ with data $(G_{1},\dotsc,G_{r};L_{1},\dotsc,L_{s})$ we mean all extensions $L/K$ such that the following two conditions hold.
   • Write $\tilde{L}/K$ for the Galois closure of $L/K$. Then $\operatorname{Gal}(\tilde{L}/K)$ isomorphic to one of $G_{1},\dotsc,G_{s}$.
@@
 Then $\mathscr{X}(\mathcal{O}_{L})_{\mathrm{new}}=\emptyset$ for 100% of $L\in\mathcal{L}$, when ordered by norms of discriminants.
-If $X=C$ is a smooth projective absolutely irreducible curve over a number field $K$, then
+If $X=C$ is a smooth projective absolutely irreducible curve over a number field $K$, and $L/K$ is a finite extension, then
 $\mathscr{X}(\mathcal{O}_{L})_{\mathrm{new}}\;=\;C(L)_{\mathrm{new}}\;\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{P\in C(L)\;:K(P)=L\}.$
@@
 Let $X/\mathbb{Q}$ be a hyperbolic curve such that $\mathscr{X}(\mathbb{Z})=\emptyset$, but $\mathscr{X}(\mathbb{Z}[(1+\sqrt{5})/2])\neq\emptyset$.
-A positive proportion of quartic number fields $L$, when ordered by absolute discriminant, contain $\mathbb{Q}(\sqrt{5})$ ([4, Corollary 1.2]) and therefore $\mathscr{X}(\mathcal{O}_{L})\neq\mathscr{X}(\mathbb{Z})$ for those fields.
+A positive proportion of quartic number fields $L$, when ordered by absolute discriminant, contain $\mathbb{Q}(\sqrt{5})$ ([5, Corollary 1.2]) and therefore $\mathscr{X}(\mathcal{O}_{L})\neq\mathscr{X}(\mathbb{Z})$ for those fields.
 We mention a few special cases of Conjecture 2 in the literature.
@@
 Siksek [30] shows that $\mathscr{X}(\mathcal{O}_{L})=\emptyset$ for $100\%$ of $L\in\mathcal{L}$, subject to a mild restriction on the mod $\ell$ representation of $J_{C}$.
-The purpose of this paper is to prove Conjecture 2 for several hyperbolic curves, with $\mathcal{L}$ the set of all cubic fields. All these hyperbolic curves have infinitely many cubic points.
+The purpose of this paper is to prove Conjecture 2 for several hyperbolic curves, with $\mathcal{L}$ the set of all quadratic or cubic fields.
+Theorem 4.
+Let $g\in\mathbb{Q}[T]$ be squarefree, of even degree $d\geq 6$, and write $G=\operatorname{Gal}(g)$ for the Galois group of $g$.
+Suppose there is an element $\sigma\in G$ acting freely on the roots of $g$.
+Let $C/\mathbb{Q}$ be the hyperelliptic curve defined by $Y^{2}=g(T)$, and write $J$ for its Jacobian.
+Suppose
+  1. (a)
+    either $d\geq 10$;
+  2. (b)
+    or $d=6$ and $J(\mathbb{Q})$ is finite;
+  3. (c)
+    or $d=8$ and $J(\mathbb{Q})$ is finite or $J$ is simple.
+Then $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of quadratic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
+We note that if $g$ is irreducible, then there is an element $\sigma\in G$ acting freely on its roots [17].
+Naturally, as $C$ is hyperelliptic, it has infinitely many quadratic points, but the above theorem is saying that the quadratic fields over which they appear are relatively rare.
+A famous theorem of Ogg [27] asserts that there are $19$ values of $N$ for which which $X_{0}(N)$ is hyperelliptic:
+  • genus $2$: $N=22$, $23$, $26$, $28$, $29$, $31$, $37$, $50$;
+  • genus $3$: $N=30$, $33$, $35$, $39$, $40$, $41$, $48$;
+  • genus $4$: $N=47$;
+  • genus $5$: $N=46$, $59$;
+  • genus $6$: $N=71$.
+Corollary 5.
+Let $N\neq 37$ be any of the above values for which $X_{0}(N)$ is hyperelliptic.
+Then $X_{0}(N)(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of quadratic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
+Theorem 6.
+Let $F(T,Y)\in\mathbb{Q}[T,Y]$ have degree $3$ in $Y$, and suppose $F$ is irreducible over $\overline{\mathbb{Q}}$.
+Let $C/\mathbb{Q}$ be the smooth projective curve with plane affine model $F(T,Y)=0$.
+Suppose $\operatorname{genus}(C)\geq 8$.
+Let $\Delta_{Y}(F)\in\mathbb{Q}[T]$ be the discriminant of $F$ with respect to $Y$, and write this as
+$\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
+where $g$, $h\in\mathbb{Q}[T]$ with $g$ squarefree.
+Suppose that $g$ has even degree, and that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$.
+Then $C(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of cubic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
+We note that the curve $C$ in Theorem 6 has infinitely many cubic points. Theorems 7 and  8 concern cubic points on certain hyperbolic curves, all of which have infinitely many cubic points.
 Theorem 7.
@@
 Then $X(L)_{\mathrm{new}}=\emptyset$ for $100\%$ of cubic fields $L/\mathbb{Q}$, when ordered by absolute discriminant.
-To motivate what comes next, we mention the following “single source theorem” due to the authors [14, Theorem 1.1].
-In this context, by a primitive degree $n$ point on $C/K$, we mean an algebraic point $P\in C(\overline{K})$ such that $K(P)/K$ has degree $n$, and the Galois group of its Galois closure is a primitive subgroup of $S_{n}$.
-Theorem.
-Let $K$ be a number field. Let $C/K$ be a curve of genus $g$, and write $J$ for the Jacobian of $C$. Let $n\geq 2$ and suppose
-(3) $\begin{cases}g>(n-1)^{2}&\text{if $n\geq 3$}\\ g\geq 3&\text{if $n=2$}.\end{cases}$
-Suppose that $A(K)$ is finite for every abelian subvariety $A/K$ of $J$ of dimension $\leq n/2$.
-If $C$ has infinitely many primitive points of degree $n$, then there is a degree $n$ morphism $\varphi:C\rightarrow\mathbb{P}^{1}$ defined over $K$ such that all but finitely many primitive degree $n$ divisors are fibres $\varphi^{*}(t)$ with $t\in\mathbb{P}^{1}(K)$.
-This theorem is called the “Single Source Theorem” since, with finitely many exceptions, all primitive degree $n$ points come from a single source which is the morphism $\varphi:C\rightarrow\mathbb{P}^{1}$.
-In view of our Conjecture 2, it makes sense to consider the family of degree $n$ number fields we obtain from the fibres of this morphism, and to attempt to show that this family contains $0\%$ of degree $n$ number fields.
-We shall give a result in this spirit.
-Let $C$ be a smooth projective absolutely irreducible curve over $K$, and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a morphism of degree $n\geq 2$.
-Then the curve $C$ possesses a plane affine (possibly singular) model
-C:F(T,Y)=0C\;:\;F(T,Y)=0
-where $F(T,Y)\in\mathbb{Q}(T)[Y]$ has degree $n$ in $Y$, is irreducible over $\overline{\mathbb{Q}}(T)$, and where the morphism $\varphi$ is given by the map $(T,Y)\mapsto T$.
-Hilbert’s irreducibility theorem asserts the existence of a thin set $\Theta\subset\mathbb{Q}$, such that for $t\in\mathbb{Q}-\Theta$, the polynomial $F(t,Y)\in\mathbb{Q}[Y]$ is irreducible of degree $n$.
-Thus $C$ has infinitely many degree $n$ points.
-For $t\in\mathbb{Q}-\Theta$ we write $K_{t}$ for the degree $n$ number field defined by $F(t,Y)=0$, and let $\Delta(\mathcal{O}_{K_{t}})$ be the discriminant of the ring of integers $\mathcal{O}_{K_{t}}$ of $K_{t}$.
-We define the discriminant set of $F$ to be
-$\operatorname{Disc}_{F}\;=\;\{\lvert\Delta(\mathcal{O}_{K_{t}})\rvert\;:\;t\in\mathbb{Q}-\Theta\}.$
-We shall give a mild sufficient condition for the set $\operatorname{Disc}_{F}$ to have natural density 0 among the positive integers.
-Whilst this is a long way off from Conjecture 2, it does harmonise with it.
+The four modular curves in Theorem 8 do not satisfy the hypothesis of Theorem 6. Nevertheless, the result follows from the proof of Theorem 6 together with further computations.
+Theorems 67 and 8 make use of results due to Bhargava, Taniguchi and Thorne, recalled in Section 5, which count (by absolute discriminant) cubic number fields satisfying local restrictions.
+For number fields of general degree, there are no such results yet.
+However, in certain settings we are able to say something about the discriminants of the primitive degree $n$ number fields for which there are new points.
+Recall that a degree $n$ number field $L/\mathbb{Q}$ is primitive if its Galois group is a primitive subgroup of $S_{n}$; equivalently, $L/\mathbb{Q}$ is primitive if the only subfields $\mathbb{Q}\subseteq K\subseteq L$ are $K=\mathbb{Q}$ and $K=L$.
 Theorem 9.
-Let $\Delta_{Y}(F)\in\mathbb{Q}(T)$ be the discriminant of $F$ with respect to $Y$, and write this as
+Let $n\geq 2$.
+Let $F(T,Y)\in\mathbb{Q}[T,Y]$ have degree $n$ in $Y$, and suppose $F$ is irreducible over $\overline{\mathbb{Q}}$.
+Let $C/\mathbb{Q}$ be the smooth projective curve with plane affine model $F(T,Y)=0$.
+Suppose
+(3) $\operatorname{genus}(C)\;>\;n^{2}-n+1.$
+Let $\Delta_{Y}(F)\in\mathbb{Q}[T]$ be the discriminant of $F$ with respect to $Y$, and write this as
 $\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
-where $g\in\mathbb{Q}[T]$ is a squarefree polynomial and $h(T)\in\mathbb{Q}(T)$ is a rational function.
-Suppose $g$ has even degree.
-Let $G=\operatorname{Gal}(g)$.
-Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$.
-Then there is some $0<\alpha<1$ such that, as $X\rightarrow\infty$,
-$(\operatorname{Disc}_{F}\cap[1,X])\;=\;O\left(\frac{X}{(\log{X})^{\alpha}}\right).$
-In particular, the discriminant set $\operatorname{Disc}_{F}$ has density 0 among the natural numbers.
+where $g$, $h\in\mathbb{Q}[T]$ with $g$ squarefree.
+Suppose that $g$ has even degree, and that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$.
+Let $\mathcal{L}$ be the set of all degree $n$ primitive number fields $L/\mathbb{Q}$.
+Let
+$\operatorname{Disc}(C,\mathcal{L})=\{\lvert\Delta(\mathcal{O}_{L})\rvert\;:\;\text{$L\in\mathcal{L}$ and $C(L)_{\mathrm{new}}\neq\emptyset$}\}.$
+Then $\operatorname{Disc}(C,\mathcal{L})$ has density 0 among the natural numbers. More precisely, there is some $0<\alpha<1$ such that
+$(\operatorname{Disc}(C,\mathcal{L})\cap[1,X])\;=\;O\left(\frac{X}{(\log{X})^{\alpha}}\right)$
+as $X\rightarrow\infty$.
+The paper is organized as follows.
+Section 2 counts, using a theorem of Delange, integers where the prime factors of the squarefree part belongs to certain special subsets of primes.
+In Section 3 we consider $F(T,Y)\in\mathbb{Q}[T,Y]$ of degree $n$ in $Y$ and, under certain hypotheses, we give an upper bound for the number of discriminants of degree $n$ number fields arising from polynomials $F(t,Y)$ with $t\in\mathbb{Q}$.
+We apply this in Section 4, together with a theorem of Vojta, to deduce Theorem 9 and Theorem 4.
+In Section 5 we recount theorems due to Bhargava, Taniguchi and Thorne that count cubic number fields with certain restrictions on ramified primes; these together with techniques introduced earlier in the paper allow us to prove Theorem 6.
+The proofs of Theorems 7 and  8 are given in Sections 6 and  7 respectively, and make use of results proved in Section 5 together with explicit computations.
+Finally, in Section 8 we suggest research projects, based on the ideas of this paper, that some readers might be interested to pursue.
 We are grateful to Maarten Derickx and Victor Flynn for helpful discussions.
 2. Counting integers with certain restricted factorisations
-Let $\mathbb{P}$ be the set of prime numbers and let $\mathcal{P}\subseteq\mathbb{P}$. Following Serre [21], we call $\mathcal{P}$ regular of density $\alpha>0$ if
+Let $\mathbb{P}$ be the set of prime numbers and let $\mathcal{P}\subseteq\mathbb{P}$. Following Serre [29], we call $\mathcal{P}$ regular of density $\alpha>0$ if, for $\operatorname{Re}(s)>1$,
 (4) $\sum_{p\in\mathcal{P}}\frac{1}{p^{s}}\;=\;\alpha\cdot\log\left(\frac{1}{s-1}\right)\;+\;\theta_{A}(s)$
@@
   • for every sufficiently large prime $p$, we have $p\in\mathcal{P}$ if and only if $\sigma_{p}\in\mathscr{C}$ where $\sigma_{p}$ is a Frobenius element of $G$ corresponding to $p$.
-By the Chebotarev Density Theorem [21, Proposition 1.5], if $\mathcal{P}$ is Frobenian of density $\alpha>0$ then it is regular of density $\alpha>0$.
+By the Chebotarev Density Theorem [29, Proposition 1.5], if $\mathcal{P}$ is Frobenian of density $\alpha>0$ then it is regular of density $\alpha>0$.
 Theorem 10.
@@
 $a_{n}=\begin{cases}1&\text{if $\operatorname{ord}_{p}(n)$ is even for all $p\in S$}\\ 0&\text{otherwise}.\end{cases}$
-Then $N_{S}(X)=\sum_{n\leq X}a_{n}$. We shall make use of a Tauberian theorem due to Delange [23, page 350] to estimate $N_{S}(X)$.
+Then $N_{S}(X)=\sum_{n\leq X}a_{n}$. We shall make use of a Tauberian theorem due to Delange [32, page 350] to estimate $N_{S}(X)$.
 Consider the Dirichlet series
@@
 where $\phi$ is holomorphic at $s=1$. Hence
-$D(s)\;=\;\frac{g_{0}(s)}{(s-1)^{1+\omega}}.$
+$D(s)\;=\;\frac{g_{0}(s)}{(s-1)^{1+\omega}}$
 where $g_{0}(s)=\exp(\phi(s))$ is holomorphic and non-vanishing at $s=1$, and $\omega=-\alpha$; the notation here is chosen to match that of Theorem II.7.28 of [32].
@@
 This gives the theorem with $C_{S}=\exp(\phi(1))/\Gamma(1-\alpha)$. ∎
-3. Proof of Theorem 6
+3. Degree $n$ points from a single source
+To motivate what comes next, we mention the following “single source theorem” due to the authors [19, Theorem 1.1].
+In this context, by a primitive degree $n$ point on $C/K$, we mean an algebraic point $P\in C(\overline{K})$ such that $K(P)/K$ is a primitive extension of degree $n$.
+Theorem.
+Let $K$ be a number field. Let $C/K$ be a curve of genus $g$, and write $J$ for the Jacobian of $C$. Let $n\geq 2$ and suppose
+(5) $\begin{cases}g>(n-1)^{2}&\text{if $n\geq 3$}\\ g\geq 3&\text{if $n=2$}.\end{cases}$
+Suppose that $A(K)$ is finite for every abelian subvariety $A/K$ of $J$ of dimension $\leq n/2$.
+If $C$ has infinitely many primitive points of degree $n$, then there is a degree $n$ morphism $\varphi:C\rightarrow\mathbb{P}^{1}$ defined over $K$ such that all but finitely many primitive degree $n$ divisors are fibres $\varphi^{*}(t)$ with $t\in\mathbb{P}^{1}(K)$.
+This theorem is called the “Single Source Theorem” since, with finitely many exceptions, all primitive degree $n$ points come from a single source which is the morphism $\varphi:C\rightarrow\mathbb{P}^{1}$.
+A variant of this theorem is due to Derickx [8].
+In view of our Conjecture 3, it makes sense to consider the family of degree $n$ number fields we obtain from the fibres of this morphism, and to attempt to show that this family contains $0\%$ of degree $n$ number fields.
+We shall give a result in this spirit.
+Let $C$ be a smooth projective absolutely irreducible curve over $\mathbb{Q}$, and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a morphism of degree $n\geq 2$.
+Then the curve $C$ possesses a plane affine (possibly singular) model
+C:F(T,Y)=0C\;:\;F(T,Y)=0
+where $F(T,Y)\in\mathbb{Q}[T,Y]$ has degree $n$ in $Y$, is irreducible over $\overline{\mathbb{Q}}$, and where the morphism $\varphi$ is given by the map $(T,Y)\mapsto T$.
+Let $\Theta\subset\mathbb{Q}$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ is either reducible, or has degree <n<n.
+Hilbert’s irreducibility theorem asserts that $\Theta$ is contained in a thin set.
+Thus $C$ has infinitely many degree $n$ points.
+For $t\in\mathbb{Q}-\Theta$ we write $L_{t}$ for the degree $n$ number field defined by $F(t,Y)=0$, and let $\Delta(\mathcal{O}_{L_{t}})$ be the discriminant of the ring of integers $\mathcal{O}_{L_{t}}$ of $L_{t}$.
+We define the discriminant set of $F$ to be
+$\operatorname{Disc}_{F}\;=\;\{\lvert\Delta(\mathcal{O}_{L_{t}})\rvert\;:\;t\in\mathbb{Q}-\Theta\}.$
+We shall give a mild sufficient condition for the set $\operatorname{Disc}_{F}$ to have natural density 0 among the positive integers.
+Whilst this is a long way off from Conjecture 3, it does harmonise with it.
+Theorem 11.
+Let $\Delta_{Y}(F)\in\mathbb{Q}(T)$ be the discriminant of $F$ with respect to $Y$, and write this as
+$\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
+where $g$, $h\in\mathbb{Q}[T]$, with $g$ squarefree.
+Suppose $g$ has even degree.
+Let $G=\operatorname{Gal}(g)$.
+Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$.
+Then there is some $0<\alpha<1$ such that, as $X\rightarrow\infty$,
+$(\operatorname{Disc}_{F}\cap[1,X])\;=\;O\left(\frac{X}{(\log{X})^{\alpha}}\right).$
+In particular, the discriminant set $\operatorname{Disc}_{F}$ has density 0 among the natural numbers.
 Lemma 12.
 Suppose $F$ satisfies the hypotheses of Theorem 11.
-Then there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even.
+Then there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))$ is even.
 Proof.
 Let $t\in\mathbb{Q}-\Theta$, and write $t=a/b$ where $a$, $b$ are coprime integers. We note that
-$\Delta(\mathcal{O}_{K_{t}})/\Delta_{Y}(F(t,Y))\;\in\;(\mathbb{Q}^{\times})^{2}.$
+$\Delta(\mathcal{O}_{L_{t}})/\Delta_{Y}(F(t,Y))\;\in\;(\mathbb{Q}^{\times})^{2}.$
 Thus
-$\Delta(\mathcal{O}_{K_{t}})/g(t)\;\in\;(\mathbb{Q}^{\times})^{2}.$
+$\Delta(\mathcal{O}_{L_{t}})/g(t)\;\in\;(\mathbb{Q}^{\times})^{2}.$
 As $g$ has even degree, say $m$, we can write
@@
 where $G$ is homogeneous, and so
-$\Delta(\mathcal{O}_{K_{t}})/G(a,b)\;\in\;(\mathbb{Q}^{\times})^{2}.$
+$\Delta(\mathcal{O}_{L_{t}})/G(a,b)\;\in\;(\mathbb{Q}^{\times})^{2}.$
 Thus, for any prime $p$,
-$\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))\;\equiv\;\operatorname{ord}_{p}(G(a,b))\pmod{2}$.
+$\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))\;\equiv\;\operatorname{ord}_{p}(G(a,b))\pmod{2}$.
 By assumption, there is some $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$.
 Therefore, by the Chebotarev Density Theorem, there is a set of primes $S$, having positive regular density, such that the congruence $G(u,v)\equiv 0\pmod{p}$ does not have solutions with coprime $u$, $v$.
-Hence for $p\in S$, we have $\operatorname{ord}_{p}(G(a,b))=0$, so $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even.
+Hence for $p\in S$, we have $\operatorname{ord}_{p}(G(a,b))=0$, so $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))$ is even.
 ∎
@@
 Theorem 11 follows immediately from Theorem 10 and Lemma 12. ∎
-4. Counting low degree fields with local restrictions
-In this section we recount some well-known theorems which count number fields of degrees $2$, 33, 44, 55, with certain local restrictions.
-We shall not need the most general form of these theorems, and content ourselves with stating simplified special cases.
-The first concerns quadratic fields, and is a special case of [7, Proposition A.1].
-Theorem 9 (Ellenberg, Pierce and Wood).
-Let $T$ be a finite set of primes. Write $N_{T}^{\pm}(X)$ for the number of isomorphism classes of quadratic fields $K$ satisfying
-  • $0<\pm\Delta_{K}\leq X$;
-  • all $p\in T$ are unramified in $K$.
-Then
-$N_{T}^{\pm}(X)\;=\;C_{T}\cdot X+O_{T}(X^{1/2})$
-where $C=1/2\zeta(2)$ and
-$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{1}{p+1}\right).$
-Most importantly for us, we shall need a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications.
+4. Proofs of Theorems 4 and 9
+We shall need the following theorem of Vojta [36, Corollary 0.3].
+Theorem (Vojta).
+Let $C$ be a curve defined over a number field $K$. Let $n$ be a positive integer and let $\varphi:C\rightarrow\mathbb{P}^{1}$ be a non-constant morphism defined over $K$. Suppose
+(6) $\operatorname{genus}(C)-1\;>\;(n-1)\cdot\deg(\varphi).$
+Then the set
+$\{P\in C(\overline{K})\;:\;\text{$[K(P):K]\leq n$ and $K(\varphi(P))=K(P)$}\}$
+is finite.
+Proof of Theorem 9.
+Let $C/\mathbb{Q}$ and $F(T,Y)\in\mathbb{Q}[T,Y]$ and $\mathcal{L}$ be as in the statement of Theorem 9.
+In particular $F$ has degree $n$ in $Y$.
+Let $\varphi:C\rightarrow\mathbb{P}^{1}$ be defined on the affine patch $F(T,Y)=0$ by $(T,Y)\mapsto T$; this has degree $n$.
+Let $L\in\mathcal{L}$.
+Thus $L$ is a primitive degree $n$ number field.
+Let $P\in C(L)_{\mathrm{new}}$.
+Then $\mathbb{Q}(P)=L$.
+Moreover, $\mathbb{Q}(\varphi(P))\subseteq L$, and therefore, as $L$ is primitive, $\mathbb{Q}(\varphi(P))=\mathbb{Q}$ or $L$.
+We note that assumption (3) in Theorem 9 implies (6).
+It follows from Vojta’s theorem that for all but finitely many $L\in\mathcal{L}$, any $(T,X)\in C(L)_{\mathrm{new}}$ satisfies $T\in\mathbb{Q}$.
+Thus, with at most finitely many exceptions, the elements of $\operatorname{Disc}(C,\mathcal{L})$ belongs to $\operatorname{Disc}_{F}$.
+Theorem 9 follows immediately from Theorem 11.
+∎
+Proof of Theorem 4.
+Let $F(T,Y)=Y^{2}-g(T)$; the discriminant with respect to $Y$ is $\Delta_{Y}(F)=2^{2}g(T)$.
+Let $d=\deg(g)$.
+Suppose first that $d\geq 10$.
+Thus $C$ has genus $\geq 4$, and so (3) is satisfied.
+Theorem 4 follows immediately from Theorem 9.
+We now consider $d=6$ and $d=8$ (and thus $C$ has genus $2$ or $3$).
+We suppose that assumptions (b), (c) in the statement of Theorem 4 hold.
+We invoke [18, Proposition 18], which is a standard consequence of the proof by Faltings [10] of the Bombieri-Lang conjecture for subvarieties of abelian varieties.
+We conclude the existence of finitely many effective degree $2$ divisors $D_{1},\dots,D_{m}$ such that
+$C^{(2)}(\mathbb{Q})=\bigcup_{i=1}^{m}\lvert D_{i}\rvert$
+where $\lvert D\rvert$ denotes the complete linear series corresponding to $D$.
+An easy application of the Riemann-Roch theorem and Clifford’s theorem [16, Theorem IV.5.4] shows that $\lvert D\rvert=\{D\}$ unless $D$ is a hyperelliptic divisor, in which case $\lvert D\rvert=\{\varphi^{*}(t):t\in\mathbb{P}^{1}(\mathbb{Q})\}$.
+Thus
+$\{P\in C(\overline{\mathbb{Q}})\;:\;\text{$[\mathbb{Q}(P):\mathbb{Q}]=2$ and $\mathbb{Q}(\varphi(P))=\mathbb{Q}(P)$}\}$
+is finite. The theorem now follows from the proof of Theorem 9. ∎
+Proof of Corollary 5.
+It is known, e.g.
+[4], that $J_{0}(N)(\mathbb{Q})$ is finite for all NN such that $X_{0}(N)$ is hyperellitic, except for $N=37$.
+The hyperelliptic polynomials $g$ for these hyperelliptic modular curves are available in the computer algebra package Magma [2].
+We computed the Galois groups and checked that the hypotheses of Theorem 4 apply.
+∎
+Remark.
+A recent paper of Najman and Trbović [25] studies primes that ramify in the field of definition of non-exceptional non-cuspidal quadratic points on hyperelliptic modular curves $X_{0}(N)$ with $N\neq 37$.
+It is possible to deduce Corollary 5 from their results, together with Theorem 10.
+5. Proof of Theorem 6
+We shall need a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications.
 The following two corollaries are easy consequences of that theorem.
 Corollary 13.
-Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $K$ satisfying
+Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $L$ satisfying
   1. (i)
-    $\lvert\Delta_{K}\rvert\leq X$;
+    $\lvert\Delta_{L}\rvert\leq X$;
   2. (ii)
-    $\operatorname{ord}_{p}(\Delta_{K})=0$ for all $p\in T$.
+    $\operatorname{ord}_{p}(\Delta_{L})=0$ for all $p\in T$.
 Then
@@
 Corollary 14.
-Let $T$ be a finite set of primes, with $3\notin T$. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $K$ satisfying
+Let $T$ be a finite set of primes, with $2$, $3\notin T$. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $L$ satisfying
   1. (i)
-    $\lvert\Delta_{K}\rvert\leq X$;
+    $\lvert\Delta_{L}\rvert\leq X$;
   2. (ii)
-    $\operatorname{ord}_{p}(\Delta_{K})$ is even for all $p\in T$.
+    $\operatorname{ord}_{p}(\Delta_{L})$ is even for all $p\in T$.
 Then
@@
 $C_{T}\;=\;\frac{1}{3\zeta(3)}\cdot\prod_{p\in T}\left(1-\frac{p}{p^{2}+p+1}\right).$
-Proof of Corollaries 10 and 11.
-Theorem 1.3 of [1] expresses the density $C_{T}^{\pm}$ as a product of local densities, which depend on the splitting type of $p$ in $K$ (either totally split, partially split, inert, partially ramified, totally ramified).
+Proof of Corollaries 13 and 14.
+Theorem 1.3 of [1] expresses the density $C_{T}$ as a product of local densities, which depend on the splitting type of $p$ in $L$ (either totally split, partially split, inert, partially ramified, totally ramified).
 In Corollary 13 the allowed splitting types for $p\in T$ are totally split, partially split, inert, and the corollary follows immediately from the aforementioned [1, Theorem 1.3].
-Next, we note that for $p\neq 3$, the valuation $\operatorname{ord}_{p}(\Delta_{K})$ is either 0, 11 or $2$ according to whether $p$ is unramified in $K$, partially ramified in $K$, or totally ramified in $K$.
-Thus, condition (ii) is equivalent to saying that for every $p\in T$ the splitting type of $p$ is either totally split, partially split, inert or totally ramified.
+We turn to the proof of Corollary 14.
+For this we note that for $p\neq 2$, $3$, the valuation $\operatorname{ord}_{p}(\Delta_{L})$ is either 0, 11 or $2$ according to whether $p$ is unramified in $L$, partially ramified in $L$, or totally ramified in $L$; this follows immediately from Lemma 15 below.
+Thus, condition (ii) of Corollary 14 is equivalent to saying that for every $p\in T$ the splitting type of $p$ is either totally split, partially split, inert or totally ramified.
 The result again follows from [1, Theorem 1.3].
 ∎
-The following is a special case of [7, Theorem 4.1].
-Theorem 12 (Ellenberg, Pierce and Wood).
-Let $T$ be a finite set of primes. Write $\tilde{N}_{T}(X)$ for the number of isomorphism classes of non-$D_{4}$ quartic fields $K$ satisfying
-  • $0<\lvert\Delta_{K}\rvert\leq X$;
-  • all $p\in T$ are unramified in $K$.
-Then
-$\tilde{N}_{T}(X)\;=\;C_{T}\cdot X+O_{T,\epsilon}(X^{23/24+\epsilon})$
-where
-$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{p^{2}+2p+1}{p^{3}+p^{2}+2p+1}\right).$
-and $C>0$ is some fixed absolute constant.
-Cohen, Diaz y Diaz and Olivier [4] show that the number of $D_{4}$-extensions with absolute discriminant at most $X$ is $\tilde{c}\cdot X$ where $c\approx 0.052326$, but there does not seem to be a count for $D_{4}$-extension with local restrictions.
-Finally, the following is a special case of [7, Theorem 5.1].
-Theorem 13 (Ellenberg, Pierce and Wood).
-Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of quintic fields $K$ satisfying
-  • $0<\lvert\Delta_{K}\rvert\leq X$;
-  • all $p\in T$ are unramified in $K$.
-Then
-$N_{T}(X)\;=\;C_{T}\cdot X+O_{T,\epsilon}(X^{199/200+\epsilon})$
-where
-$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{p^{3}+2p^{2}+2p+1}{p^{4}+p^{3}+2p^{2}+2p+1}\right).$
-and $C>0$ is some fixed absolute constant.
+The following result is well-known. We include the proof for the convenience of the reader.
+Lemma 15.
+Let $L$ be a number field of degree $n$. Let $p>n$ be a rational prime and write
+$p\mathcal{O}_{L}=\mathfrak{P}_{1}^{e_{1}}\cdots\mathfrak{P}_{r}^{e_{r}}$
+where the $\mathfrak{P}_{i}$ are distinct prime ideals. Let $f_{i}$ be the inertial degree of $\mathfrak{P}_{i}$. Then
+$\operatorname{ord}_{p}(\Delta_{L})=\sum_{i=1}^{r}f_{i}(e_{i}-1).$
+Proof.
+The condition $p>n$ ensures that $p\nmid e_{i}$, so the primes $\mathfrak{P}_{i}$ are tamely ramified in $L$.
+Write $\mathcal{D}_{L}$ for the different ideal of $L$.
+Then $\operatorname{ord}_{\mathfrak{P}_{i}}(\mathcal{D}_{L})=e_{i}-1$ (Theorem III.2.6 of [26]).
+The lemma follows as $\operatorname{Norm}(\mathcal{D}_{L})=\lvert\Delta_{L}\rvert$ (Theorem III.2.9 of [26]).
+∎
 Lemma 16.
@@
 However, as $\sigma\rightarrow 1^{+}$,
-$\frac{\sum_{p\in R}\frac{1}{p^{\sigma}}}{-\log(\sigma-1)}\;\leq\;\frac{\sum_{p\in R}\frac{1}{p}}{-\log(\sigma-1)}\rightarrow 0$
-since $\sum_{p\in\mathbb{P}}1/p^{\sigma}=-\log(\sigma-1)+O(1)$. This contradicts the assumption that $\delta>0$. ∎
+$\frac{\sum_{p\in R}\frac{1}{p^{\sigma}}}{-\log(\sigma-1)}\;\leq\;\frac{\sum_{p\in R}\frac{1}{p}}{-\log(\sigma-1)}\rightarrow 0.$
+This contradicts the assumption that $\delta>0$. ∎
 Corollary 17.
-Let $S$ be a set of primes having regular density <1<1.
-  1. (a)
-    Let $d=2$, 33 or $5$. Then the proportion of degree $d$ number fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
-  2. (b)
-    Let $d=4$. Then the proportion of non-$D_{4}$ quartic fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
+Let $S$ be a set of primes having regular density <1<1. Then the proportion of cubic number fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
 Proof.
-Write $\mathcal{F}(X)$ for the set of quintic fields with absolute discriminant at most $X$, and write $\mathcal{F}_{S}(X)$ for the subset whose discriminants are supported on $S$.
-To prove the corollary for $d=5$, we want to show that
+Write $\mathcal{F}(X)$ for the set of cubic fields with absolute discriminant at most $X$, and write $\mathcal{F}_{S}(X)$ for the subset whose discriminants are supported on $S$. We want to show that
 (7) $\lim_{X\rightarrow\infty}\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}=0.$
@@
 Hence
-$0\leq\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}\leq\frac{N_{p_{1},\dotsc,p_{k}}(X)}{N_{\emptyset}(X)}\rightarrow\prod_{i=1}^{k}\left(1-\frac{p_{i}^{3}+2p_{i}^{2}+2p_{i}+1}{p_{i}^{4}+p_{i}^{3}+2p_{i}^{2}+2p_{i}+1}\right)$
-as $X\rightarrow\infty$, where the limit follows from Theorem 13.
+$0\leq\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}\leq\frac{N_{p_{1},\dotsc,p_{k}}(X)}{N_{\emptyset}(X)}\rightarrow\prod_{i=1}^{k}\left(1-\frac{p_{i}+1}{p_{i}^{2}+p_{i}+1}\right)$
+as $X\rightarrow\infty$, where the limit follows from Corollary 13.
 From Lemma 16, by choosing $p_{1},\dotsc,p_{k}$ be a suitable (finite) subset of $R$, we can ensure that the product on the right hand-side is as small as we like.
-This completes the proof of (a) for $d=5$.
-The other cases are similar.
+This completes the proof.
 ∎
@@
 Corollary 19.
-Let $F(T,Y)\in\mathbb{Q}(T)[Y]$ has degree $3$ in $Y$, and be irreducible over $\mathbb{Q}(T)$, and $\Theta$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ either has degree <n<n or is reducible (this set is necessarily thin by Hilbert’s irreducibility theorem).
-Let $\Delta_{Y}(F)\in\mathbb{Q}(T)$ be the discriminant of $F$ with respect to $Y$, and write this as
+Let $F(T,Y)\in\mathbb{Q}[T,Y]$ be irreducible and have degree $3$ in $Y$.
+Let $\Theta$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ either has degree <3<3 or is reducible.
+Let $\Delta_{Y}(F)\in\mathbb{Q}[T]$ be the discriminant of $F$ with respect to $Y$, and write this as
 $\Delta_{Y}(F)\;=\;g(T)\cdot h(T)^{2}$
-where $g\in\mathbb{Q}[T]$ is a squarefree polynomial and $h(T)\in\mathbb{Q}(T)$ is a rational function.
+where $g$, $h\in\mathbb{Q}[T]$ with $g$ squarefree.
 Suppose $g$ has even degree.
@@
 Proof.
-By Lemma 8, there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even, for any $K_{t}\in\mathcal{F}^{\prime}$.
+By Lemma 12, there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))$ is even, for any $L_{t}\in\mathcal{F}^{\prime}$.
 The desired result follows from Corollary 18.
+∎
+Proof of Theorem 6.
+We apply the proof of Theorem 9 with $n=3$.
+We deduce that all but finitely many cubic points on $C$ arise as fibres of the map $C\rightarrow\mathbb{P}^{1}$ given on the affine model $F(T,Y)=0$ by $(T,Y)\mapsto T$; the assumption $\operatorname{genus}(C)\geq 8$ ensures that $C$ satisfies condition (3).
+Theorem 6 follows immediately from Corollary 19.
 ∎
@@
 Nagell [22], [23], has determined all exceptional number fields where the unit rank is 0 or $1$.
-Nagell finds [18, Section 2] that the only exceptional quadratic fields are $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-3})$ which contain exceptional units $(1+\sqrt{5})/2$ and $(1-\sqrt{-3})/2$ respectively, and the only exceptional complex cubic fields are the ones with discriminants 23-23 and 31-31.
-He also showed [18, Sections 3–5] that the only exceptional real cubic fields are of the form $\mathbb{Q}(\varepsilon)$ where $\varepsilon$ is a root of
+Nagell finds [23, Section 2] that the only exceptional quadratic fields are $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-3})$ which contain exceptional units $(1+\sqrt{5})/2$ and $(1+\sqrt{-3})/2$ respectively, and the only exceptional complex cubic fields are the ones with discriminants 23-23 and 31-31.
+He also showed [23, Sections 3–5] that the only exceptional real cubic fields are of the form $\mathbb{Q}(\varepsilon)$ where $\varepsilon$ is a root of
 $f_{t}(X)=X^{3}+(t-1)X^{2}-tX-1,\qquad t\in\mathbb{Z},\quad t\geq 3$
@@
 $F(t)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\Delta_{X}(f_{t})=t^{4}+6t^{3}+7t^{2}-6t-31.$
-It follows that if a prime $p$ divides the discriminant of $\mathcal{O}_{\mathbb{Q}(\varepsilon)}$ then it divides $t^{4}+6t^{3}+7t^{2}-6t-31$ and hence the polynomial $F(X)$ has a root modulo $p$.
+It follows that if a prime $p$ divides the discriminant of $\mathcal{O}_{\mathbb{Q}(\varepsilon)}$ then it divides $F(t)$ and hence the polynomial $F$ has a root modulo $p$.
 Write $S$ for the set of primes $p$ such that $F$ has a root modulo $p$.
@@
 Write $\mathcal{F}_{S}$ for the set of cubic fields with discriminants supported on $S$.
-By the above, the exceptional fields with Galois group $S_{3}$ all belong to $\mathcal{F}_{S}$.
+By the above, the real exceptional fields with Galois group $S_{3}$ all belong to $\mathcal{F}_{S}$.
 The latter has density 0 by Corollary 17.
 This completes the proof.
+∎
 7. Proof of Theorem 8
@@
 Using Magma [2] we computed the Riemann–Roch space $L(D_{a,b,c})$ for the $32$ possible combinations $(a,b,c)$.
-For all but $4$ of these, the Riemann–Roch dimension is $1$, and therefore $D=D_{a,b,c}$.
+For all but $4$ of these, the Riemann–Roch dimension is $1$, and therefore there is precisely one effective $D$ linearly equivalent to $D_{a,b,c}$.
 For the remaining $4$, namely $(a,b,c)=(0,0,0)$, $(0,1,0)$, $(0,1,1)$, $(1,2,1)$, the Riemann–Roch dimension is $2$, and therefore there are infinitely many effective degree $3$ divisors linearly equivalent to $D_{a,b,c}$.
 Consider the case $(a,b,c)=(0,0,0)$.
-Then $D_{(0,0,0)}=3P_{1}$, and a basis for $L(D_{(0,0,0)})$ is given by $1$, $\varphi=y/x$, where these are considered as elements of the function field $\mathbb{Q}(X_{0}(64))$.
+Then $D_{(0,0,0)}=3P_{1}$, and a basis for $L(D_{(0,0,0)})$ is given by $1$, $\varphi\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}y/x$, where these are considered as elements of the function field $\mathbb{Q}(X_{0}(64))$.
 Thus
@@
 for the Magma code supporting the proof of this theorem.
+8. Suggested Future Projects
+We invite the reader to help gather further evidence towards our Conjecture 2. Here are some projects that we believe are ‘low-hanging fruit’:
+  1. (I)
+    Theorem 9 gives an upper bound for the number of discriminants of degree $n$ primitive number fields coming from degree $n$ new points on a curve $C$.
+It assumed that $C$ has a plane model of the form $F(T,Y)=0$ where $F$ has degree $n$ in $Y$.
+This assumption is equivalent to the existence of a degree $n$ morphism $\psi:C\rightarrow\mathbb{P}^{1}$.
+We expect that it should be possible to prove a similar theorem where there is a degree $n$ morphism $\psi:C\rightarrow E$ to an elliptic curve $E$.
+Here, instead of Vojta’s Theorem (recounted in Section 4), it would be useful to apply a more general theorem due to Song and Tucker [31, Proposition 2.3].
+  2. (II)
+    The proof of Theorem 6 makes use of a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications.
+Similar theorems for degrees $4$ and $5$ have been established by Ellenberg, Pierce and Wood [9, Theorems 4.1 and 5.1].
+As far as we can tell, to prove the analogue of Theorem 6 for degrees $4$ and $5$ would require a refinement of the aforementioned theorems of Ellenberg, Pierce and Wood, that consider densities of different possible ramification types for primes.
+  3. (III)
+    Corollary 5 and Theorem 8 establish Conjecture 3 in degrees $2$, 33 for certain $X_{0}(N)$.
+It should be possible to establish many similar results for other modular curves appearing in the $\mathrm{L}$-Functions and Modular Forms database [33].